# Active Preference Learning for Ordering Items In- and Out-of-sample

### v1

@@ Abstract
Learning an ordering of items based on noisy pairwise comparisons is useful when item-specific labels are difficult to assign, for example, when annotators have to make subjective assessments. Algorithms have been proposed for actively sampling comparisons of items to minimize the number of annotations necessary for learning an accurate ordering. However, many ignore shared structure between items, treating them as unrelated, limiting sample efficiency and precluding generalization to new items. In this work, we study active learning with pairwise preference feedback for ordering items with contextual attributes, both in- and out-of-sample. We give an upper bound on the expected ordering error incurred by active learning strategies under a logistic preference model, in terms of the aleatoric and epistemic uncertainty in comparisons, and propose two algorithms designed to greedily minimize this bound. We evaluate these algorithms in two realistic image ordering tasks, including one with comparisons made by human annotators, and demonstrate superior sample efficiency compared to non-contextual ranking approaches and active preference learning baselines.
@@ 1 Introduction
The success of supervised learning is built on annotating items at great volumes with small error. For subjective assessments, however, assigning a value from an arbitrary rating scale can be difficult, causing many to favor preference feedback from pairwise comparisons  (Phelps et al., 2015; Yannakakis and Martínez, 2015; Christiano et al., 2017; Zhu et al., 2023). Preference feedback is sufficient to learn an ordering of items (Fürnkranz and Hüllermeier, 2003), but for $n$ items, there are $\mathcal{O}(n^{2})$ possible pairs of items to compare. A common solution is to use crowd-sourcing  (Chen et al., 2013; Yang et al., 2021; Larkin et al., 2022), but what if the task requires certain expertise and annotations are expensive to collect? This is the case in the field of medical imaging, where annotations require trained radiologists (Phelps et al., 2015; Jang et al., 2022; Lidén et al., 2024, Appendix A). To be cost-efficient, we must select comparisons carefully, to minimize the annotations needed to learn a good ordering.
There is a large literature on ordering a set of items $\mathcal{I}$ using only noisy pairwise comparisons of actively chosen item pairs (Herbrich et al., 2006; Maystre and Grossglauser, 2017; Heckel et al., 2018; Lou et al., 2022). While robust against noise and more efficient than exhaustive comparison, these algorithms ignore any context attributes that the items may have. Say that a radiologist wants to quantify the expression of a disease in a collection of X-ray images $\mathcal{I}$. Ranking these images according to their severity $y_{i}$ using purely preference-based algorithms ignores the content $x_{i}$ of the X-rays, sampling images $i,j\in\mathcal{I}$ for comparison based only on the current estimated severity and previous feedback. However, the content of the images reveal similarities between X-rays which can inform an ordering strategy.
In applications, the set $\mathcal{I}$ we want to order is often larger than the set of items we observe during training—we may want to rank new images in relation to previous ones. However, methods based only on preference feedback cannot generalize what they learned to new items. As an alternative, active learning algorithms for classification can be used for ordering by fitting a map from pairs of contexts $x_{i},x_{j}$ to the comparison $i\succ_{?}j$, that can be applied to new items (Houlsby et al., 2011; Kirsch et al., 2019; Bengs et al., 2022). However, to the best of our knowledge, there is no analysis that justifies using these algorithms to learn a complete ordering or gives guarantees for their performance.
@@
In this work, we study active preference learning algorithms for ordering sets of items based on noisy pairwise comparisons. We propose the use of a contextual logistic preference model to support efficient ordering in-sample as well as generalization to new items. Our analysis yields the first bound on the expected ordering error achievable under this model, provided a fixed budget of comparisons (Section 4). This result justifies an active sampling principle based on the epistemic and aleatoric uncertainty in the outcome of a comparison between two items. We use this in one greedy, deterministic, and one Bayesian, stochastic algorithm (Section 5) applied to both a fully contextual preference model, and a hybrid variant with per-item parameters, aimed at overcoming potential model misspecification (Section 5.3). We evaluate these models and sampling strategies in two realistic image ordering tasks, one where the preference feedback is simulated using a logistic model, and one where the comparisons have been performed by human annotators (Section 6). We show that our sampling strategy compares favorably to active preference learning and bandit baselines, and that our hybrid model benefits both from the low variance of contextual models and the low bias of fitting per-item parameters. This manifests in faster convergence in-sample, better generalization to new items, and more efficient tuning when continuing learning on new items.
2 Ordering with active preference learning
Our goal is to learn an ordering of items $\mathcal{I}$ according to an unobserved score $y_{i}\in\mathbb{R}$, defined for each item $i\in\mathcal{I}$. The ground-truth ordering of $\mathcal{I}$ is determined by a comparison function $\pi_{ij}\coloneqq\mathds{1}[y_{i}>y_{j}]$ such that $\pi_{ij}=1$ indicates that $i$ is ranked higher than $j$. We assume there are no ties.
@@ 2 Ordering with active preference learning
where $n=|\mathcal{I}|$. This error is equivalent to the normalized Kendall’s Tau distance (Kendall, 1948). Our analysis may be readily generalized to non-uniform distributions.
We learn $h$ from preference feedback—noisy pairwise comparisons $C_{ij}\in\{0,1\}$ for items $(i,j)$ related to their score, for example, provided by one or more human annotators. $C_{ij}=1$ indicates that the annotator perceived that item $i$ has a higher score than $j$, i.e., that they prefer $i$ over $j$. Our goal is to minimize the ordering error $R_{\mathcal{I}}(h)$ for a fixed budget $T\geq 1$ of adaptively chosen comparisons.
We are particularly interested in algorithms that are guaranteed to generalize out-of-sample to unseen items. In this setting, only a subset of $\mathcal{I}_{D}\subseteq\mathcal{I}$ is available during learning, and the learner must generalize to $\mathcal{I}\setminus\mathcal{I}_{D}$. Ordering algorithms based only on preference feedback cannot solve this problem since observed comparisons are uninformative of new items. To overcome this, we let each item $i\in\mathcal{I}$ be endowed with item-specific attributes $x_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$. The active preference learning scenario proceeds as follows:
  1. 1.
    A learner is given an annotation budget $T$, a pool of items $\mathcal{I}_{D}\subseteq\mathcal{I}$ and per-item attributes $x_{i}$ for $i\in\mathcal{I}_{D}$.
  2. 2.
    Over rounds $t=1,...,T$, the learner requests a comparison of two items $i_{t},j_{t}\in\mathcal{I}_{D}$ according to a sampling criterion and receives noisy binary preference feedback $c_{t}\sim p(C_{ij})$, independently of previous comparisons.
  3. 3.
    After $T$ rounds, the learner returns an estimated comparison function $h:\mathcal{I}\times\mathcal{I}\rightarrow\{0,1\}$.
We denote the history of accumulated observations (pairs of items and comparisons) up until and including time $t$ by $D_{t}=((i_{1},j_{1},c_{1}),...,(i_{t},j_{t},c_{t}))$. To learn a good model by time $T$, an active learner must use all information in $D_{t}$ to inform its sampling criterion. A natural baseline is the uniform criterion, which selects any pair $i\neq~{}j\in\mathcal{I}_{D}$ with equal probability, ignorant of $D_{t}$. Uniform sampling gives equal weight to item pairs with similar scores and pairs with vastly different scores, which may result in many comparisons having minimal impact on the final estimate $h$.
Logistic preference feedback. We analyze the case where comparisons $C_{ij}$ are generated according to a logistic model applied to the difference between item scores, $p(C_{ij}=1)=\sigma(y_{i}-y_{j})~{},$ where $\sigma(z)=1/(1+e^{-z})$ is the logistic function and $\dot{\sigma}(z)$ its derivative at $z$. Note that $\dot{\sigma}(z)$ is also the variance of a Bernoulli variable parameterized by $\sigma(z)$.
We are particularly interested in the case where $y_{i}$ is a linear function of item attributes, $y_{i}=\theta_{*}^{\top}x_{i}~{},$ with $\theta_{*}\in\mathbb{R}^{d}$ the ground-truth coefficients. In this case, comparisons are determined by a logistic regression model applied to the difference vector between item attributes $z_{ij}\coloneqq x_{i}-x_{j}$,
@@
This model is a special case of the family of parametric models considered in Heckel et al. (2018), and is equivalent to the logistic model often used in learning to rank (Burges et al., 2005). It satisfies linear (and thus strong) stochastic transitivity (Oliveira et al., 2018). A similar score function is assumed also in Chu and Ghahramani (2005).
@@ 3 Related work
Preference learning (Fürnkranz and Hüllermeier, 2003; Chu and Ghahramani, 2005) is related to the problem of learning to rank (Burges et al., 2005; Jamieson and Nowak, 2011; Busse et al., 2012). Solving preference learning using as few, adaptively chosen comparisons as possible may be posed as an active learning (Brinker, 2004; Long et al., 2010; Silva et al., 2014; Ling et al., 2020) or bandit problem. Non-contextual active learners, such as TrueSkill (Herbrich et al., 2006; Minka et al., 2018), Hamming-LUCB (Heckel et al., 2018), and Probe-Rank (Lou et al., 2022) produce in-sample orderings, but must be updated if new items are to be ranked. This is mitigated by contextual algorithms that exploit structure between items, such as BALD (Houlsby et al., 2011). BALD, when applied to preference learning, samples items by maximizing the expected information gain, a commonly used criterion in Bayesian optimal design (Lindley, 1956). Kirsch and Gal (2022) show that many recently proposed active learning strategies may be unified in a framework based on Fisher information. Chaudhuri et al. (2015) analyze active learning for general maximum likelihood estimation, and Ash et al. (2021) extend this idea for models parameterized by deep neural networks.
@@
Bandit algorithms with relative or dueling feedback (Yue and Joachims, 2009; Bengs et al., 2021; Yan et al., 2022), have been proposed both in contextual (Dudík et al., 2015) and non-contextual settings (Yue et al., 2012). Bengs et al. (2022) proposed CoLSTIM, a contextual dueling bandit for regret minimization under linear stochastic transitivity, matching the model in (2). Di et al. (2023) gave variance-aware regret bounds for this setting. However, as argued in Appendix E, algorithms that identify the top-$k$ items as fast as possible, such as pure exploration bandits (Fang, 2022; Jun et al., 2021), can be arbitrarily bad at learning a full ordering. Related are also George and Dimitrakakis (2023) who consider the problem of finding a Kemeny ranking in non-contextual dueling bandits, and Wu et al. (2023b) who minimize the Borda regret. Our analysis uses techniques from generalized linear bandits (Filippi et al., 2010; Li et al., 2017; Faury et al., 2020; Kveton et al., 2020). For an extended review of related work, see Appendix B.
4 A bound on the ordering error
@@
We give an upper bound on the ordering error $R_{\mathcal{I}}(h)$ for a hypothesis $h$ fit using a fixed budget of comparisons, $T$. We restrict our analysis to the logistic model in (2) and denote by $R(\theta)\equiv R_{\mathcal{I}}(h_{\theta})$ the risk of the hypothesis defined by $h_{\theta}(i,j)=\mathds{1}[\theta^{\top}z_{ij}>0]$. Let $\theta_{t}$ be the maximum-likelihood estimate (MLE) of the true parameter $\theta_{*}$, defined in (9) in the Appendix, fit to $t$ rounds of feedback, $D_{t}$.
Fix $i,j\in\mathcal{I}$ and recall that $z_{ij}=x_{i}-x_{j}$. For convenience, we define $z_{t}\equiv z_{i_{t}j_{t}}$ as the difference between attributes for the pair of items selected at round $t$. Now, let $\Delta_{ij}>0$ be a lower bound on the margin of comparison such that $\left|\sigma(z_{ij}^{\top}\theta_{*})-1/2\right|>\Delta_{ij}$ for all $i,j$. Further, define $\Delta_{*}=\min_{ij}\Delta_{ij}$. Let $\mathbf{H}_{t}(\theta):=\sum_{s=1}^{t}\dot{\sigma}(z_{s}^{\top}\theta)z_{s}z_{s}^{\top}$ be the Hessian of the negative log-likelihood of observations at time $t$ under (2), given the parameter $\theta$. We define
$\tilde{\mathbf{H}}_{t}(\theta):=\sum_{i,j}w_{t}(i,j)\dot{\sigma}(z_{i,j}^{\top}\theta)z_{i,j}z_{i,j}^{\top}$
with $w_{t}(i,j))=N_{t}(i,j)/t$ where $N_{t}(i,j)$ denotes the number of times the pair $(i,j)$ has been queried up to time $t$. Hence, $\mathbf{H}_{t}(\theta_{t})=t\tilde{\mathbf{H}}_{t}(\theta_{t})$. We further note that $\mathbf{H}_{t}(\theta)$ is the observed Fisher information, and define the norm $||x||_{V}$, for a square matrix $V$, as $\|x\|_{V}=\sqrt{x^{\top}Vx}$.
We make the following assumptions for our analysis:
@@
$\mathbf{H}_{T}(\theta_{T})$ and $\mathbf{H}_{T}(\theta_{*})$ have full rank and their minimum eigenvalues are larger than some $\lambda_{0}>0$.
Assumptions 1 and 2 are required for our analysis but the learner does not have to know these bounds. Assumption 3 is naturally satisfied, for sufficiently large $T$, by most reasonable sampling strategies, or can be ensured by allowing for a burn-in phase of $O(d)$ uniformly drawn samples in the beginning of an adaptive strategy.
We start by stating the following concentration result for the deviation of $\sigma(z_{ij}^{\top}\theta_{T})$ from the true probability $\sigma(z_{ij}^{\top}\theta_{*})$. The proof of Lemma 1 is found in Appendix D and builds on results for optimistic algorithms in logistic multi-armed bandits (Filippi et al., 2010; Faury et al., 2020).
@@
$\alpha_{ij}(\Delta)$ $:=\exp\Bigg{(}\frac{-\Delta^{2}T}{8dC_{1}\left(\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T}))}\right)^{2}}\Bigg{)},\;\;\beta_{ij}(\Delta)$ $:=\exp\Bigg{(}\frac{-\Delta T}{dC_{1}\left(\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}\right)^{2}}\Bigg{)}~{}.$
@@
Here $C_{1}$ is a constant depending on $S,\lambda_{0},Q$ from Assumptions 13 (see Appendix D for definition).
The concentration result in Lemma 1 is verifiable (given by observables) since the upper bound depends only on the maximum likelihood estimate $\theta_{T}$ at time $T$, not on $\theta_{*}$. We present a sharper, unverifiable bound in Appendix D which instead depends on $\theta_{*}$ but does not suffer from the explicit scaling with $d$ in the definitions of $\alpha$ and $\beta$. The bound in Lemma 1 can also be expressed in terms of $\mathbf{H}^{-1}_{T}(\theta_{T})$ by using the equality $||z_{ij}||^{2}_{\mathbf{H}^{-1}_{T}(\theta_{T})}=\frac{1}{T}||z_{ij}||^{2}_{\hat{\mathbf{H}}^{-1}_{T}(\theta_{T})}$
As long as our sampling strategy ensures that the minimum eigenvalue of $\tilde{\mathbf{H}}_{t}(\theta_{t})$ does not tend to zero, i.e., the strategy is strongly consistent (Chen et al., 1999), we have $\alpha_{ij}(\Delta_{ij})\sim\exp\left(-\Delta_{ij}^{2}T/(\dot{\sigma}(z_{ij}^{\top}\theta_{T})^{2}\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})})\right)$ and $\beta_{ij}(\Delta_{ij})\sim\exp\left(-\Delta_{ij}T/\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}\right)$. Since $\Delta^{2}_{ij}<\Delta_{ij}<1/2$ by definition, we can view $\alpha$ as the first-order term and $\beta$ as the second-order term of our bound. Lemma 1 formally captures the intuition that it should be easier to sort when annotations contain little noise, i.e., $\dot{\sigma}(z_{ij}^{\top}\theta_{T})$ is small. Especially, we observe $\dot{\sigma}(z_{ij}^{\top}\theta_{T})\approx 0$ for pairs where $\Delta_{ij}$ is sufficiently large, causing the first-order term to vanish, leaving us with the faster decaying second-order term $\beta$.
Lemma 1 also tells us that the hardest pairs to guarantee a correct ordering for are the ones with both high aleatoric uncertainty under the MLE model, e.g., where annotators disagree or labels are noisy, captured by $\dot{\sigma}(z_{ij}^{\top}\theta_{T})$, as well as high epistemic uncertainty captured by $||z_{ij}||_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}$.
A direct consequence of Lemma 1 is the following bound,
@@
The RHS in the above inequality can be bounded further by utilizing that $\Delta_{ij}\geq|i-j|\Delta_{*}$. Together with Markov’s inequality, this yields the following bound on $P\left(R(\theta_{T})\geq\epsilon\right)$.

### v2

@@ Abstract
Learning an ordering of items based on pairwise comparisons is useful when items are difficult to rate consistently on an absolute scale, for example, when annotators have to make subjective assessments. When exhaustive comparison is infeasible, actively sampling item pairs can reduce the number of annotations necessary for learning an accurate ordering. However, many algorithms ignore shared structure between items, limiting their sample efficiency and precluding generalization to new items. It is also common to disregard how noise in comparisons varies between item pairs, despite it being informative of item similarity. In this work, we study active preference learning for ordering items with contextual attributes, both in- and out-of-sample. We give an upper bound on the expected ordering error of a logistic preference model as a function of which items have been compared. Next, we propose an active learning strategy that samples items to minimize this bound by accounting for aleatoric and epistemic uncertainty in comparisons. We evaluate the resulting algorithm, and a variant aimed at reducing model misspecification, in multiple realistic ordering tasks with comparisons made by human annotators. Our results demonstrate superior sample efficiency and generalization compared to non-contextual ranking approaches and active preference learning baselines.
@@ 1 Introduction
The success of supervised learning is built on annotating items at great volumes with small error. For subjective assessments, however, assigning a value from an arbitrary rating scale can be difficult and prone to inconsistencies, causing many to favor preference feedback from pairwise comparisons  (Yannakakis and Martínez, 2015; Christiano et al., 2017; Ouyang et al., 2022; Zhu et al., 2023). Preference feedback is sufficient to learn an ordering of items (Fürnkranz and Hüllermeier, 2003), but for $n$ items, there are $\mathcal{O}(n^{2})$ possible pairs of items to compare. A common solution is to use crowd-sourcing  (Chen et al., 2013; Yang et al., 2021; Larkin et al., 2022), but many tasks require domain expertise, making annotations expensive to collect. This is the case in the field of medical imaging, where annotations require trained radiologists (Phelps et al., 2015; Jang et al., 2022; Lidén et al., 2024; Tärnåsen and Bergström, 2023). So, how can we learn the best ordering possible from a limited number of comparisons?
Classically, this problem is solved by active learning, sampling comparisons based on preference feedback and estimated item scores (Herbrich et al., 2006; Maystre and Grossglauser, 2017; Heckel et al., 2018). However, consider a radiologist who wants to quantify the expression of a disease in a collection of X-ray images. Purely preference-based algorithms utilize only the outcomes of comparisons but ignore the contents of the X-rays, which can reveal similarities between items and inform an ordering strategy. Moreover, the set we want to order is often larger than the set of items observed during training—we may want to rank new X-rays in relation to previous ones. This cannot be solved by learning per-item scores alone. As an alternative, active learning for classification can be used to fit a map from pairs of item contexts $x_{i},x_{j}$ (e.g., the contents of images) to the comparison $i\succ_{?}j$, that can be applied to old and new items alike (Houlsby et al., 2011; Qian et al., 2015). However, as we show in Section 4, learning this map to recover a complete ordering is distinct from the tasks preference learning is commonly used for, and existing algorithms lack theoretical justification for this application. Moreover, formal results for related problems, such as contextual bandits or reinforcement learning (Das et al., 2024; Filippi et al., 2010; Zhu et al., 2023; Bengs et al., 2022), do not translate directly to effective active sampling criteria for ordering. There is a small body of work on learning a contextual model to recover the complete ordering (Jamieson and Nowak, 2011; Ailon, 2011) but these either assume noiseless preference feedback or that the noise is unrelated to the similarity of items, which is unrealistic for subjective assessments.
@@
We propose using a contextual logistic preference model to support efficient in-sample ordering and generalization to new items. Our analysis yields the first bound on the expected ordering error achievable given a collected set of comparisons (Section 4). This result justifies an active sampling principle that accounts for both epistemic and aleatoric uncertainty which we implement in a greedy deterministic algorithm called GURO (Section 5). We further propose a hybrid variant of the contextual preference model, compatible with GURO as well as existing sampling strategies, that overcomes model misspecification by adding per-item parameters (Section 5.1). We evaluate GURO and baseline algorithms in four diverse ordering tasks, three of which utilize comparisons performed by human annotators (Section 6). Our sampling strategy compares favorably to active preference learning baselines, and our hybrid model benefits both GURO and other sampling criteria, achieving the low variance of contextual models and the low bias of fitting per-item parameters. This results in faster convergence in-sample, better generalization to new items, and efficient continual learning when new items are added.
2 Ordering items with active preference learning
Our goal is to learn an ordering of items $\mathcal{I}$ according to an unobserved score $y_{i}\in\mathbb{R}$, defined for each item $i\in\mathcal{I}$. The ground-truth ordering of $\mathcal{I}$ is determined by a comparison function $\pi_{ij}\coloneqq\mathds{1}[y_{i}>y_{j}]$, where $\pi_{ij}=1$ indicates that $i$ ranks higher than $j$. We assume there are no ties.
@@ 2 Ordering items with active preference learning
where $n=|\mathcal{I}|$. This error is equivalent to the normalized Kendall’s Tau distance (Kendall, 1948).
Hypotheses $h$ are learned from preference feedback—noisy pairwise comparisons $C_{ij}\in\{0,1\}$ for items $(i,j)$ related to their score, for example, provided by human annotators. $C_{ij}=1$ indicates that an annotator perceived that item $i$ has a higher score than $j$, i.e., that they prefer $i$ over $j$. Our goal is to minimize the ordering error $R_{\mathcal{I}}(h)$ for a fixed budget $T\geq 1$ of adaptively chosen comparisons.
We are interested in contextual problems, where each item $i\in\mathcal{I}$ is endowed with item-specific attributes $x_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$. As we will see, this permits more sample-efficient ordering and learning algorithms that can order items out-of-sample, trained on comparisons of a subset of items $\mathcal{I}_{D}\subseteq\mathcal{I}$ and generalizing to $\mathcal{I}\setminus\mathcal{I}_{D}$. Ordering algorithms based only on preference feedback cannot solve this problem since observed comparisons are uninformative of new items.
Our active preference learning scenario proceeds as follows: 1) A learner is given an annotation budget $T$, a pool of items $\mathcal{I}_{D}\subseteq\mathcal{I}$ and item attributes $x_{i}$ for $i\in\mathcal{I}_{D}$. 2) Over rounds $t=1,...,T$, the learner requests a comparison of two items $i_{t},j_{t}\in\mathcal{I}_{D}$ according to a sampling criterion and receives noisy binary preference feedback $c_{t}\sim p(C_{ij})$, independently of previous comparisons. 3) After $T$ rounds, the learner returns a comparison function $h:\mathcal{I}\times\mathcal{I}\rightarrow\{0,1\}$. We denote the history of accumulated observations until and including time $t$ by $D_{t}=((i_{1},j_{1},c_{1}),...,(i_{t},j_{t},c_{t}))$.
We assume that comparisons $C_{ij}$ follow a logistic model applied to the difference between item scores, $p(C_{ij}=1)=\sigma(y_{i}-y_{j}),$ the so-called Bradley-Terry model (Bradley and Terry, 1952), which assumes linear stochastic transitivity (Oliveira et al., 2018). Throughout, $\sigma(z)=1/(1+e^{-z})$ and $\dot{\sigma}(z)$ its derivative at $z$. Specifically, we study the case where $y_{i}$ is a linear function of item attributes, $y_{i}=\theta_{*}^{\top}x_{i}~{},$ with $\theta_{*}\in\mathbb{R}^{d}$ the ground-truth coefficients. Thus, comparisons are determined by a logistic regression model applied to the attribute difference vector $z_{ij}\coloneqq x_{i}-x_{j}$,
@@
We face two kinds of uncertainty when actively learning the model in (2): epistemic and aleatoric. Epistemic uncertainty, or model uncertainty, is the uncertainty about the true parameter $\theta_{*}$, while aleatoric uncertainty is the irreducible uncertainty about labels due to noisy annotation.
@@ 3 Related work
Preference learning (Fürnkranz and Hüllermeier, 2003; Chu and Ghahramani, 2005) is related to the problem of learning to rank (Burges et al., 2005; Busse et al., 2012). When using adaptively chosen comparisons it may be posed as an active learning or bandit problem (Brinker, 2004; Long et al., 2010; Silva et al., 2014; Ling et al., 2020). Non-contextual active learners, such as TrueSkill (Herbrich et al., 2006; Minka et al., 2018), Hamming-LUCB (Heckel et al., 2018), and Probe-Rank (Lou et al., 2022) produce in-sample preference orderings, but must be updated if new items are to be ranked. Contextual algorithms, such as BALD (Houlsby et al., 2011), mitigate this by exploiting item structure and Kirsch and Gal (2022) show that many recently proposed contextual active learning strategies may be unified in a framework based on Fisher information. Similarly, methods have been proposed to recover a linear preference model by adaptively sampling paired comparisons (Qian et al., 2015; Massimino and Davenport, 2021; Canal et al., 2019). Still, this setting differs from ours in that we emphasize recovering the full ordering, not perfectly estimating the parameters. While it is true that knowing the parameters is sufficient to order the list, reducing uncertainty for all parameters equally will likely be wasteful (see Section 4). Ailon (2011) offer guarantees for ordering using contextual features in the noiseless setting, while Jamieson and Nowak (2011) analyze the setting where noise is unrelated to item similarity.
Bandits:
@@
Bandit algorithms with relative or dueling feedback (Yue and Joachims, 2009; Bengs et al., 2021; Yan et al., 2022) also learn from pairwise comparisons, and have been proposed both in contextual (Dudík et al., 2015) and non-contextual settings (Yue et al., 2012) to minimize regret or identify top-$k$ items. Bengs et al. (2022) proposed CoLSTIM, a contextual dueling bandit for regret minimization under linear stochastic transitivity, matching (2), and Di et al. (2023) gave variance-aware regret bounds for this setting. However, algorithms that find the top-$k$ items, such as pure exploration bandits (Fang, 2022; Jun et al., 2021), can be arbitrarily bad at learning a full ordering (see Appendix D). Related are also George and Dimitrakakis (2023) who learn Kemeny rankings in non-contextual dueling bandits, and Wu et al. (2023b) who minimize Borda regret. Zhu et al. (2023) studies the problem of estimating a preference model from offline data. Our analysis uses techniques from logistic bandits (Filippi et al., 2010; Li et al., 2017; Faury et al., 2020; Kveton et al., 2020).
RLHF:
Preference learning is commonly used when training large language models through reinforcement learning with human feedback (RLHF) (Christiano et al., 2017; Bai et al., 2022; Ouyang et al., 2022; Wu et al., 2023a). In this line of work, Zhu et al. (2023) provide guarantees on the sample complexity of learning a preference model from offline data. They leverage similar tools from statistical learning and bandits as we do. In contrast to their work, we provide sampling strategies for the online setting. Mehta et al. (2023) consider active learning for RLHF in a dueling bandit framework where the goal is to optimize a contextual version of the Borda regret. Concurrent work by Mukherjee et al. (2024) and Das et al. (2024) studies a similar problem, as we do here, in the RLHF setting but with the objective to identify an optimal policy in a contextual bandit with dueling feedback. In contrast to their objective, we are interested in recovering the ordering of items. Das et al. (2024) use similar bandit techniques as we do, and their selection criterion, when adapted for ordering, corresponds to our NormMin baseline (see Section 6).
4 Which comparisons result in a good ordering?
We give an upper bound on the ordering error $R_{\mathcal{I}}(h)$ for a hypothesis $h$ fit using a set of $T$ comparisons. The bound is retrospective, attempting to answer the question “if we collect comparisons $D_{T}$, how good is our resulting model at ordering the items in $\mathcal{I}$?”. In Section 5, we use insights from the result to design an active learning algorithm.
@@ 4 Which comparisons result in a good ordering?
We restrict our analysis to the logistic model in (2) and denote by $R(\theta)\equiv R_{\mathcal{I}}(h_{\theta})$ the risk of the hypothesis defined by $h_{\theta}(i,j)=\mathds{1}[\theta^{\top}z_{ij}>0]$. Recall that $z_{ij}=x_{i}-x_{j}$ for $i,j\in\mathcal{I}$, and define $z_{t}\equiv z_{i_{t}j_{t}}$ as the difference between attributes for the pair of items selected at round $t$. Let $\theta_{t}$ be the maximum-likelihood estimate (MLE) fit to $t$ rounds of feedback, $D_{t}$
$\theta_{t}=\operatorname*{arg\,max}_{\theta}\sum_{s=1}^{t}\left(c_{s}\log\sigma(\theta^{\top}z_{s})+(1-c_{s})(1-\sigma(\theta^{\top}z_{s}))\right)~{}.$ (3)
Let $\Delta_{ij}>0$ lower bound the margin of comparison, $|\sigma(z_{ij}^{\top}\theta_{*})-1/2|>\Delta_{ij}$ for all $i,j\in\mathcal{I}$ and define $\Delta_{*}=\min_{i\neq j}\Delta_{ij}/|i-j|$. Next, let $\mathbf{H}_{t}(\theta):=\sum_{s=1}^{t}\dot{\sigma}(z_{s}^{\top}\theta)z_{s}z_{s}^{\top}$ be the Hessian of the negative log-likelihood of observations at time $t$ under (2), given the parameter $\theta$, also known as observed Fisher information. We define $\tilde{\mathbf{H}}_{t}(\theta):=\frac{1}{t}\mathbf{H}_{t}(\theta)$. For a square matrix $V$, we define $\|x\|_{V}=\sqrt{x^{\top}Vx}$. We make the following assumptions for our analysis:
@@
$\mathbf{H}_{T}(\theta_{T})$ and $\mathbf{H}_{T}(\theta_{*})$ have full rank and minimum eigenvalues larger than $\lambda_{0}>0$.
Assumption 1 implies that $\theta_{*}$ lies in some ball with radius $S$ and cannot have unbounded coefficients. Assumption 2 states that there exists an upper bound on the norm of the feature vectors. This assumption is trivially satisfied whenever we have a finite set of data points. Both assumptions 1 and 2 are standard in the bandit literature and only required for our analysis. Assumption 3 is naturally satisfied for sufficiently large $T$ by any sampling strategy with support on $d$ linearly independent vectors, or can be ensured by allowing for a burn-in phase of $d$ samples in the beginning of an adaptive strategy. Assumption 3 ensures the uniqueness of $\theta_{t}$.
We start by stating the following concentration result for the deviation of $\sigma(z_{ij}^{\top}\theta_{T})$ from the true probability $\sigma(z_{ij}^{\top}\theta_{*})$. The proof of Lemma 1 is found in Appendix C and builds on results for optimistic algorithms in logistic multi-armed bandits (Filippi et al., 2010; Faury et al., 2020).
@@
$\alpha_{ij}(\Delta):=\exp\Bigg{(}\frac{-\Delta^{2}T}{8dC_{1}(\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})})^{2}}\Bigg{)},\;\;\;\beta_{ij}(\Delta):=\exp\Bigg{(}\frac{-\Delta T}{dC_{1}\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}}\Bigg{)}~{}.$
@@
$C_{1}$ depends on $S,\lambda_{0},Q$ from Assumptions 13 (see Appendix C for definition and proof).
The concentration result in Lemma 1 is verifiable (given by observables) since the upper bound depends only on the maximum likelihood estimate $\theta_{T}$ at time $T$, not on $\theta_{*}$. We present a sharper, unverifiable bound in Appendix C which instead depends on $\theta_{*}$ but does not suffer from the explicit scaling with $d$ in the definitions of $\alpha$ and $\beta$. The bound in Lemma 1 can also be expressed in terms of $\mathbf{H}^{-1}_{T}(\theta_{T})$ by using the equality $||z_{ij}||^{2}_{\mathbf{H}^{-1}_{T}(\theta_{T})}=\frac{1}{T}||z_{ij}||^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}.$ As long as our sampling strategy ensures that the minimum eigenvalue of $\tilde{\mathbf{H}}_{t}(\theta_{t})$ does not tend to zero, i.e., the strategy is strongly consistent (Chen et al., 1999), we have $\alpha_{ij}(\Delta_{ij})\sim\exp[-\Delta_{ij}^{2}T/(\dot{\sigma}(z_{ij}^{\top}\theta_{T})^{2}\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})})]$ and $\beta_{ij}(\Delta_{ij})\sim\exp[-\Delta_{ij}T/\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}]$. Since $\Delta^{2}_{ij}<\Delta_{ij}<1/2$ by definition, we can view $\alpha$ as the first-order term and $\beta$ as the second-order term of our bound.
Lemma 1 formally captures the intuition that it should be easier to sort when annotations contain little noise, i.e., $\dot{\sigma}(z_{ij}^{\top}\theta_{T})$ is small. Especially, we observe $\dot{\sigma}(z_{ij}^{\top}\theta_{T})\approx 0$ for pairs where $\Delta_{ij}$ is sufficiently large, causing the first-order term to vanish, leaving us with the faster decaying second-order term $\beta$. Lemma 1 also tells us that the hardest pairs to guarantee a correct ordering for are the ones with both high aleatoric uncertainty under the MLE model, e.g., where annotators disagree or labels are noisy, captured by $\dot{\sigma}(z_{ij}^{\top}\theta_{T})$, as well as high epistemic uncertainty captured by $||z_{ij}||_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}$.
A direct consequence of Lemma 1 is the following bound on the ordering error of $h_{\theta_{T}}$ over $\mathcal{I}$,
@@
The right-hand side in the above inequality can be bounded further by utilizing that $\Delta_{ij}\geq|i-j|\Delta_{*}$. Together with Markov’s inequality, this yields the following bound on $P(R(\theta_{T})\geq\epsilon)$.